OpenStudy (anonymous):

Determine the slope of the line that contains the points (0, 2) and (-5, 1). (Points : 1) 1/5 5/2 -1/5 5

pooja195 (pooja195):

y\[ \frac{ y2-y1 }{ x2-x1 }\]

pooja195 (pooja195):

now just plug in your coordinates

OpenStudy (anonymous):

What is the slope of the line defined by the table? -7/3 3/7 -3/7 7/3

pooja195 (pooja195):

can you do the 1st one before you post another one?

OpenStudy (anonymous):

ok

pooja195 (pooja195):

\[\frac{ 1-2 }{ -5-0 }\]

pooja195 (pooja195):

there now solve :) @iGreen can you finish up please i need to go :)

OpenStudy (igreen):

Sure thing

OpenStudy (anonymous):

1/5

OpenStudy (igreen):

Correct.

OpenStudy (igreen):

1 - 2 = -1 -5 - 0 = -5 -1/-5 = 1/5

OpenStudy (anonymous):

ok can you help me with other ?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

I might be able to, close this and open a new one.